OpenStudy (anonymous):

8x+2y=-2 y=-5x+1 solve by substitution

OpenStudy (anonymous):

Thank you for helping me !

OpenStudy (anonymous):

We know what 'y' is, so all we have to do is go back to the equation: 8x + 2y = -2 and substitute. We get: 8x + 2(-5x + 1) = -2. As you can see, I plugged in the value for 'y' (-5x + 1) into this equation. All we have to do now is solve! 8x -10x + 2 = -2 ...Used distribution. -2x + 2 = -2 ...Solving out. -2x = -4 ...Almost done! x = 2 ...Solved!!! Now we take the value for x, and use this to find y. y = -5x + 1, so: y = -5(2) + 1; that means y = -9. ^_^

OpenStudy (anonymous):

I hope I helped. If you don't understand something I did, please let me know. :)

OpenStudy (anonymous):

I understand thanks for your help!